### v1

# Under-Canopy Navigation using Aerial Lidar Maps

Abstract
Autonomous navigation in unstructured natural environments poses a significant challenge. In goal navigation tasks without prior information, the limited look-ahead of onboard sensors utilised by robots compromises path efficiency. We propose a novel approach that leverages an above-the-canopy aerial map for improved ground robot navigation. Our system utilises aerial lidar scans to create a 3D probabilistic occupancy map, uniquely incorporating the uncertainty in the aerial vehicle’s trajectory for improved accuracy. Novel path planning cost functions are introduced, combining path length with obstruction risk estimated from the probabilistic map. The D* Lite algorithm then calculates an optimal (minimum-cost) path to the goal. This system also allows for dynamic replanning upon encountering unforeseen obstacles on the ground. Extensive experiments and ablation studies in simulated and real forests demonstrate the effectiveness of our system.
Index Terms:
Navigation, Field Robots, Robotics and Automation in Agriculture and Forestry
I Introduction
Long-range navigation in natural environments such as forests remains a hard problem. Despite recent advances in perception and planning, ground robots may struggle to identify traversable areas in vegetated environments and the onboard sensor’s limited field of view can lead to inefficient paths, frequently facing dead-ends during deployment. Aerial data, e.g., lidar or satellite images, collected prior to the ground vehicle deployment, could offer a more global perspective, potentially allowing the generation of paths that avoid some of the obstacles visible from above (Fig. 1).
Figure 1: Sensors mounted on a ground robot can suffer from limited range and occlusions. This local view often precludes efficient long-distance navigation in complex scenarios, leading to dead-ends or unsafe paths (bottom figure: yellow arrows). Aerial lidar data (top left figure) acquired prior to operation can provide meaningful information for global planning (top right and bottom figures: magenta paths) allowing efficient navigation.
Although the problem of ground navigation using aerial information has been investigated previously [1, 2], leveraging airborne data for under-canopy robotic navigation poses several challenges. The different perspectives between above- and under-canopy views can lead to the occlusion of ground obstacles from aerial images or under-sampling of ground lidar points (the relevant points for ground-vehicle planning) caused by canopy cover. Moreover, the sensing noise and inevitable error in the estimated aerial vehicle’s trajectory further hinder the creation of accurate aerial maps that can guide the ground robot. Additionally, the path planner must incorporate occupancy belief encoded in the probabilistic map in addition to path length to minimise the risk of lengthy detours and, ultimately, mission failure.
We address these problems in a novel framework for under-canopy navigation in forests assisted by above-canopy aerial lidar scans. Our system builds a 3D probabilistic occupancy map from overhead lidar observations and uses occupancy probabilities to estimate the likelihood of ground-level obstruction throughout the map. Standard occupancy mapping techniques approach the problem as a pure mapping task with perfect sensor poses. Ignoring the inevitable error in aerial sensor’s pose estimates (produced by, e.g., pose-graph optimisation) can result in poor mapping accuracy, especially at longer ranges in aerial scenarios (lidar hit-point position error increases approximately linearly with the product of angle and range). To address this issue, our mapping system explicitly takes into account pose-estimation uncertainty and averages occupancy probabilities over the set of possible aerial vehicle trajectories using Monte Carlo sampling. The aerial map is then used to plan a safe and efficient path for the ground robot before its deployment. We introduce two path planning cost functions designed to strike a balance between minimising the travel distance and avoiding potential obstructions (inferred from the occupancy map). Our path planner is based on D* Lite [3] which supports efficient replanning in case the ground robot encounters a previously unseen obstacle along its current path. Finally, we evaluate the performance of our full system and its components in extensive experiments conducted in simulated and real complex forest environments. In summary, the main contributions of this work are the following:
  1. 1.
    A framework for navigation in dense forest environments using aerial lidar observations as prior information;
  2. 2.
    A mapping algorithm to estimate the obstructed areas under the canopy using airborne lidar data;
  3. 3.
    A Monte Carlo sampling method to propagate uncertainty from the aerial sensor trajectory to 3D occupancy maps;
  4. 4.
    Experimental evaluation in simulation and real-world experiments in a challenging forest area using a mobile tracked robot.
II Related Work
II-A Mapping using aerial sensing modalities
Navigation in unstructured environments has drawn attention recently [4, 5, 6], however, the utilisation of aerial sensing data to support ground navigation is a less explored idea. To create traversability maps used in global planning, Vandapel et al. [7] first segment ground points from airborne lidar, then calculate costs based on the “vegetationess” metric (i.e., ratio of ground points and total number of points falling in a 2D grid region), and a vehicle mobility analysis above the terrain. Alternatively, a combination of geometric features (e.g., ground elevations) and semantic terrain classes from airborne sensing data (lidar, color and multispectral image) is used to estimate traversability costs [8]. In a similar fashion, [9] combines semantic terrain classes of earth observation data, geometric features (e.g., slope) and also proprioceptive locomotion data (e.g., velocities, IMU) of different robotic platforms to generate traversal costs conditioned on the platform type and terrain class. Lastly, Stentz et al. [10] learn high-cost features in a self-supervised fashion by fusing air-ground sensing data (satellite image and ground-view lidar).
Despite their relevance, [7, 8, 10] neglect free-space information and [7, 8] measurement uncertainty in their costmap definitions, which are crucial elements in determining traversable areas. Our approach relies on explicit probabilistic reasoning over voxel occupancy (free, occupied, or unknown), while also accounting for sensor noise and pose-estimation error. While we focus our method on obstructions inferred from occupancy maps, complementary notions of traversability from learning methods (e.g., semantic-based costs) could also be fused in our framework.
II-B Sensor pose uncertainty in occupancy maps
Standard occupancy mapping approaches [11] and many of its variants [12, 13, 14] all assume that sensor poses are perfectly known at the time lidar measurements are acquired. However, this approximation can easily lead to high occupancy errors if localisation uncertainty increases. To address this problem, O’Callaghan et al. [15] propose to incorporate sensor and localisation uncertainty in Gaussian Process (GP) occupancy mapping by considering inputs as noisy observations drawn from a Gaussian distribution, and integrating them in the kernel function. However, such modifications increase the computational cost of the approach by an order of magnitude, which may limit its application in large datasets even in offline settings.
In this work, we revisit and extend the idea of using Monte Carlo (MC) sampling to incorporate localisation uncertainty in 2D occupancy maps. This approach is first proposed in [16] and is used more recently in [17]. The former integrates out poses from occupancy distributions using sample trajectories obtained from a particle filter, whereas the later performs the MC integration in the log-odds occupancy ratio. While we also marginalise out poses from occupancy probabilities using MC sampling, unlike [16, 17] our approach generates samples from the approximated posterior of a modern pose-graph SLAM system. Furthermore, we apply this idea to 3D mapping and conduct extensive experimental evaluations in simulated forest environments.
II-C Path planning
Graph-based (D*, A*) or sampling-based (Rapidly Exploring Random Trees - RRT*) methods have been broadly used for global planning in costmaps [18]. When applied to occupancy maps, occupancy probabilities are often thresholded to classify grid cells as obstacle or free space, subsequently assigning high and low costs to them [19, 20]. Other works [7, 21] propose heuristic functions of terrain features to build costmaps used in the referred planners. However, planning on thresholded occupancy maps or feature-based costs without uncertainty information can create a false sense of confidence in the generated path, ultimately leading to long detours.
Banfi et al. [22] address the problem of planning under uncertainty by generating multiple path hypotheses in unknown areas and reasoning on locally acquired information from a “next-best view” (NBV) pose to take the best route.
Unlike the above approaches, our system generates a minimum-cost path to the goal by minimising cost functions that blend obstruction risk with path length. Additionally, we use D* Lite to efficiently replan a new optimal path each time the ground robot faces a previously unmapped obstacle along its current path.
III System Overview
We divide the problem of forest navigation using uncertain prior maps from aerial lidar into two tasks: 1) offline mapping and global planning using airborne sensing data; and 2) online execution of waypoints and obstacle avoidance using onboard perception on a ground robot. Since this work is focused on offline mapping and planning, we utilise an off-the-shelf onboard navigation system (local navigation module [23]) for the online execution (task 2) in our experiments.
Figure 2: Overview of our proposed navigation system. First, aerial lidar data are fused with uncertain sensor poses to generate a 3D probabilistic occupancy map. Next, the 3D map is further processed to generate a 2D ground obstruction map. Finally, obstruction scores are passed through a cost function to the D* Lite planner to produce global paths that guide the ground robot during operation.
Fig. 2 summarises our proposed system. The aerial mapping (Sec. IV) starts aggregating airborne lidar scans and estimating the sensor trajectory using a localisation system, namely Wildcat SLAM [24]. Then, we estimate the ground height from the point cloud (Sec. IV-A), in addition to building a 3D probabilistic occupancy map from the aerial lidar observations (Sec. IV-B). For improved map accuracy, we also encode the uncertainty from sensor poses into occupancy distributions by means of a Monte Carlo integration method. Next, the ground height estimates and occupancy values are fed into our scoring function to estimate the obstructed (i.e., non-traversable) areas over the terrain (Sec. IV-C). Obstruction scores are finally integrated in two evaluated cost functions, expected cost and log-reachability cost, ultimately used by D* Lite to generate optimal paths.
IV Aerial Mapping
IV-A Ground Filtering
The ground filtering module estimates the ground height from 3D point clouds, producing a 2D grid map representation of the terrain’s supporting surface, where each grid square cell (side: $0.25$ m) stores the height information. Height estimates are further used in our mapping approach as references during the estimation of obstructed regions above the terrain.
We first apply the Cloth Simulation Filtering (CSF) [25] to segment ground points from the global point cloud. CSF vertically ($Z$) inverts the point cloud data and simulates gravity acting on a semi-rigid cloth – modeled as interconnected mass-spring nodes – covering the inverted surface. Then, CSF segments ground points by retaining the closest points in the cloud to the CSF nodes within a height threshold in each grid cell. For cells without ground points, we use the closest CSF node as a virtual ground point. Finally, the minimum height point in each grid cell is used as the ground height estimate. By operating on inverted point clouds, CSF is less susceptible to segmentation errors (e.g., lidar canopy reflections or high vegetation mistakenly classified as ground) and generalises well to distinct environments using only few parameters.
IV-B Occupancy mapping with imperfect pose estimates
In addition to the ground elevations, occupancy estimates of the environment are also used as inputs to the obstruction estimation module. We process lidar points above ground using the Occupancy Homogeneous Mapping (OHM) [26] to build an occupancy grid map. OHM estimates an occupancy probability for each 3D grid element (voxel) by taking into account lidar hits and pass-throughs in each voxel. Standard occupancy mapping methods such as OHM assume lidar poses are perfectly known. However, in practice, poses are estimated by other state estimation systems (e.g., pose-graph optimisation, or some other form of SLAM) which are inevitably subject to estimation error due to sensor noise. Ignoring pose estimation errors leads to highly inaccurate maps. We address this challenge by incorporating the estimation uncertainty of the aerial vehicle’s trajectory provided by the pose-estimation system into occupancy mapping.
The occupancy mapping module estimates the occupancy probability of each voxel given aerial lidar measurements. We denote the map by $\{p(m_{i}\mid\mathcal{Z})\}_{i}$ where $m_{i}$ is a Bernoulli random variable indicating the occupancy of the $i$th voxel, and $\mathcal{Z}$ denotes the set of all measurements from the aerial vehicle (e.g. lidar, IMU, and GNSS). This probability can be computed as follows:
$p(m_{i}\mid\mathcal{Z})=\int p(m_{i}\mid\mathcal{X},\mathcal{Z})\,p(\mathcal{X}\mid\mathcal{Z})\,\mathrm{d}\mathcal{X},$ (1)
where $\mathcal{X}$ denotes the robot’s trajectory (represented by a finite set of poses). The first integrand $p(m_{i}\,|\,\mathcal{X},\mathcal{Z})$ in (1) is the voxel occupancy probability computed by standard occupancy mapping systems given trajectory $\mathcal{X}$. Note that voxel occupancy $m_{i}$ is conditionally independent from IMU and GNSS measurements given the trajectory $\mathcal{X}$. Thus one can further simplify this term as $p(m_{i}\,|\,\mathcal{X},\mathcal{Z}_{\text{lidar}})$ where $\mathcal{Z}_{\text{lidar}}\subset\mathcal{Z}$ denotes the set of lidar measurements. The second integrand $p(\mathcal{X}\,|\,\mathcal{Z})$ is the posterior distribution over robot trajectory given all measurements. This term is computed by the underlying SLAM system (e.g., Wildcat’s pose-graph optimisation module [24]). We use Monte Carlo integration to approximate the intractable integral in (1). This is done by sampling $N$ trajectories $\{\mathcal{X}^{[k]}\}_{k=1}^{N}$ from the pose-graph SLAM posterior $p(\mathcal{X}\mid\mathcal{Z})$ and computing the following expression,
$p(m_{i}\mid\mathcal{Z})\approx\frac{1}{N}\sum_{k=1}^{N}p(m_{i}\mid\mathcal{X}^{[k]},\mathcal{Z}).$ (2)
Equation (2) suggests the following scheme for incorporating pose-estimation uncertainty into occupancy mapping: (i) build an occupancy map using any standard occupancy mapping method such as OHM for each trajectory sample $\mathcal{X}^{[k]}$; and (ii) for each voxel compute the average of the resulting $N$ predicted occupancy probabilities. We sample trajectories from the marginal covariances of individual pose posteriors. After solving the pose-graph SLAM optimisation, such covariances can be approximated by leveraging the relevant diagonal blocks of the information matrix representing the full posterior $p(\mathcal{X}\mid\mathcal{Z})$. Hereafter we use the term expected occupancy map to refer to the result of this module.
IV-C Obstruction score estimation
Figure 3: We reason on free, occupied and unknown voxels by integrating lidar measurements in an occupancy grid map. Later the occupancy probabilities are fused in our scoring method to estimate the final obstruction map.
We use our 3D expected occupancy map and the estimated ground height to assign a simple obstruction risk score to each ground-level 2D cell. This score is computed for each vertical column of $n$ 3D voxels starting from the ground cell as illustrated in Fig. 3. Formally, the score function $b:\mathbb{Z}\times\mathbb{Z}\to[0,1]$ assigns a value between zero and one to each 2D cell (i.e., the grid $XY$ coordinates of the corresponding column in the world frame). A larger value of $b(s)$ for a cell $s\triangleq(x,y)$ indicates a higher risk of obstruction at that cell. The number $n$ of column voxels is chosen according to the clearance height ($h$ in Fig. 3) required by the specified robot to pass through a column. Therefore, taller robots require using a larger $n$.
A cell $s$ might be considered as obstructed if at least one of the voxels in the corresponding column is occupied. This choice leads to $b(s)=1-\prod_{i\in\mathcal{I}_{s}}p(m_{i}=0\mid\mathcal{Z})$. However, because of the complexity of forest environments and the sparsity of aerial lidar rays reaching the ground level, the resulting score function will likely fail to correctly capture traversable cells where only the first voxel above ground is occupied by grass or an overhanging foliage is occupying the top voxel in the stack. An effective heuristic estimate is to use instead the weighted average of voxel occupancy probabilities within a column as the corresponding obstruction score, where the (fixed) weights reflect the significance of voxel occupancy. The resulting score function $b$ is given by
$b(s)\triangleq\sum_{i\in\mathcal{I}_{s}}\frac{w_{i}}{\sum_{i\in\mathcal{I}_{s}}w_{i}}\,p(m_{i}=1\mid\mathcal{Z}),$ (3)
where $\mathcal{I}_{s}$ denotes the set of indices of the $n$ voxels that are part of the vertical stack at cell $s$, and $p(m_{i}=1\mid\mathcal{Z})$ is the (expected) occupancy probability of voxel $i$ given the aerial measurements. Specifically, in our work $n=4$ which allows for a robot height of 111\,1m using a 0.250.250.25\,0.25m voxel size, and the voxels weights $w_{i}$ from bottom to top are set to $(1,2,2,2)$. The rationale behind these values is to assign a smaller weight to the first voxel above ground level since the tracked mobile robot used in our experiments can better navigate obstacles immediately above the ground (e.g., small rocks, grass).
Finally, note that the obstruction score $b(s)$ is computed for the corresponding voxel-sized 2D cell $s$ which is typically much smaller than the robot’s footprint. Therefore, in the definition of path costs in Sec. V, we use the (worst-case) footprint obstruction score $b_{\text{max}}(s)\triangleq\max_{s_{i}\in\mathcal{F}(s)}\,b(s_{i})$ among cells in the robot footprint $\mathcal{F}(s)$ when the robot’s center is inside cell $s$. For convenience, we use a circular footprint, avoiding the need to consider different heading angles of the robot.
V Global Path Planning
Our path planning module plans a 2D path (over the ground cells) for the ground robot given a pair of start and goal locations. A path $\mathcal{P}$ is a sequence of adjacent ground cells $\mathcal{P}=(s_{0},s_{1},\ldots,s_{\ell})$ starting from the current cell $s_{0}$ and ending at the goal cell $s_{\ell}=s_{\text{goal}}$. The cost of transitioning from cell $s$ to one of its adjacent cells $s^{\prime}$ (hereafter, the edge cost) is denoted by $c(s,s^{\prime})\in\mathbb{R}^{+}$. We use D* Lite [3] to find a minimum-cost path where the path cost is given by the sum of its edge costs,
$J({\mathcal{P}})\triangleq\sum_{i=0}^{\ell-1}c(s_{i},s_{i+1}).$ (4)
D* Lite allows us to efficiently replan an optimal path from the current cell to the goal based on the latest edge costs upon encountering a blockage. In the following, we introduce two choices used in our simulations and experiments for the edge costs $c(s,s^{\prime})$ that incorporate both the footprint obstruction score $b_{\text{max}}(s)$ (reflecting the risk of choosing a blocked path) and the travelled distance.
V-1 Expected Cost
This edge cost function is inspired by the expected traversal cost proposed in [27, Eq. 8]. Formally, the edge cost between two adjacent cells $s_{i}$ and $s_{i+1}$ in a given path is defined as follows,
$c(s_{i},s_{i+1})=b_{\text{max}}(s_{i+1})\,C_{\text{obst}}+\left(1-b_{\text{max}}(s_{i+1})\right)||s_{i+1}-s_{i}||$ (5)
where $C_{\text{obst}}\in\mathbb{R}^{+}$ is a penalty parameter aimed at discouraging the robot from entering cells that have a high obstruction score. Higher values of $C_{\text{obst}}$ result in a higher penalty for overlapping with potentially obstructed cells. Intuitively, if we interpret $b_{\text{max}}(s)\in[0,1]$ as the obstruction probability when the robot’s center is inside cell $s$, the above expression becomes the expected value of the following function:
$(s_{i},s_{i+1})\mapsto\begin{cases}C_{\text{obst}}&\text{if $s_{i+1}$ is blocked,}\\ \|s_{i+1}-s_{i}\|&\text{if $s_{i+1}$ is free.}\end{cases}$ (6)
In this way, the cost function represents an expectation of the transition cost where $C_{\text{obst}}$ is the additional detour cost incurred when overlapping an obstructed cell, and $b_{\text{max}}(s)$ is the likelihood. D* Lite also requires a heuristic cost function $h(s_{0},s_{i})$ that must be non-negative and backward consistent, i.e., it obeys $h(s_{0},s_{0})=0$ and $h(s_{0},s_{i})\leq h(s_{0},s_{i-1})+c(s_{i-1},s_{i})$. In our method, the heuristic cost is defined as
$h(s_{0},s_{i})=(1-b_{\text{min}})d(s_{0},s_{i}),$ (7)
where $d(s,s^{\prime})$ is the octile [^1] Octile distance: $d(s,s^{\prime})\triangleq\max(\Delta x,\Delta y)+(\sqrt{2}-1)\min(\Delta x,\Delta y)$, where $\Delta x=|s^{\prime}[0]-s[0]|$ and $\Delta y=|s^{\prime}[1]-s[1]|$. distance between $s$ and $s^{\prime}$ considering also diagonal movements in a grid, and $b_{\text{min}}\triangleq\min_{s}\,b(s)$ is the minimum obstruction score found in our map.
V-2 Log-Reachability cost
Our second path cost is based on the reachability metric proposed in [28]. Informally speaking, reachability of a path reflects the likelihood of reaching the goal cell without hitting any blocked cells. We approximate the product integral derived in [28, Eq. 14] for computing the reachability of the goal-directed path $\mathcal{P}$ by the following expression,
$R(\mathcal{P})\approx\prod_{i=0}^{\ell-1}\left(1-b_{\text{max}}({s_{i+1}})\right)^{||s_{i+1}-s_{i}||}.$ (8)
We find the optimal path by minimising the negative log-reachability, yielding the following edge cost,
$c(s_{i},s_{i+1})$ $=-\log\left(1-b_{\text{max}}(s_{i+1})\right)\,{||s_{i+1}-s_{i}||}.$ (9)
The heuristic cost function used for the reachability cost is
$h(s_{0},s_{i})=-\log(1-b_{\text{min}})\,d(s_{0},s_{i})$ (10)
where, as before, $d(s,s^{\prime})$ denotes the octile distance between $s$ and $s^{\prime}$, and $b_{\text{min}}$ is the minimum obstruction score across the entire map.
VI Simulations
We validated our proposed approach using the Gazebo simulator [^2] https://gazebosim.org/. We created three synthetic forests using vegetation models of the NEGS-UGV Dataset [29], and acquired simulated airborne lidar scans of these forest worlds. In the following, we first evaluate the accuracy of our proposed expected occupancy mapping approach, and then analyse the performance of our full navigation system.
VI-A Ablation Study: Expected Occupancy Mapping
The objective is to evaluate the accuracy of our expected occupancy map (with localisation uncertainty) and compare it with the standard occupancy map for different levels of error in the sensor trajectory. We created three random forest worlds named Forest I, II and III, and simulated three different aerial trajectories (one per scenario), as shown in Fig. 4, of a VLP-16 lidar sensor flying above the canopy and scanning the forests. We constructed ground-truth occupancy maps of each scenario by simulating an ideal sensor (no noise in range measurements or trajectory) and processing the ground-truth point clouds and trajectories using OHM. Then, we perturbed each sensor pose along the trajectories with independent, additive Gaussian noise $\sim\mathcal{N}\left(0,\sigma^{2}\right)$ on each component to simulate a (perturbed) localisation estimate of the sensor trajectory. For each perturbation level, specified by standard deviation values, we generated 10 perturbed trajectories and point clouds for subsequent comparisons. We also added Gaussian noise with a fixed standard deviation ($\sigma_{r}=0.01\,$m) to range measurements to simulate sensor uncertainty. For our expected occupancy mapping (Sec. IV-B) we chose $N=20$.
Figure 4: (a): Example of synthetic forest created in Gazebo where we simulated lidar sensors moving above the canopy to evaluate our method for incorporation of uncertainty from sensor poses to occupancy values. (b) to (d): Ground-truth trajectories of the airborne sensor and the corresponding registered scans for the datasets acquired in the synthetic environments Forest I, II and III, respectively. After adding perturbations to sensor poses, a perturbed point cloud of each dataset was produced to be processed by our approach.
With full knowledge of ground-truth occupancy probabilities from the ideal sensor, and the perturbed datasets, we can further inspect the maps’ accuracy using the Kullback-Leibler Divergence (KLD) of voxel occupancy probabilities between the expected and standard occupancy maps with respect to the ground-truth map.
In Forest I, we analysed the isolated effect of perturbations in sensor’s translation and orientation, respectively. Since the percentage of occupied voxels (about $10\%$) in the ground-truth map is considerably lower than the free counterparts (about $80\%$), we also computed the average KLD of expected and standard occupancy maps per voxel class (free, occupied and unknown) in this first scenario. Ground-truth probability ($p_{\text{gt}}$) thresholds are used to classify voxels as free ($p_{\text{gt}}<0.45$), unknown ($0.45\leq p_{\text{gt}}\leq 0.55$) or occupied ($p_{\text{gt}}>0.55$).
Figure 5: Comparison of the expected occupancy against the standard occupancy map for different perturbation values added to sensor poses. (a) and (b) show the average KLD values in Forest I dataset for perturbations in position only and orientation only, respectively. KLD values are depicted per voxel class on the ground-truth dataset. (c) and (d) illustrate the KLD for all voxels classes in datasets Forest II and Forest III, respectively, for combined perturbation levels (position and orientation simultaneously) where $\sigma=(0.015\text{m},0.015\text{m},0.015\text{m},0.11^{\circ},0.11^{\circ},0.11^{\circ})$. In the figures, noise refers to perturbations added to sensor poses.
Results illustrated in Fig. 5 (a) and (b) show that for both perturbations (position and orientation), the maps generated using expected occupancy presented lower average KLD for all perturbation levels, and that this behaviour was consistent among all voxels classes, confirming that our method improves the quality of maps on occupied, free and unknown regions. We also assessed the effect of combined perturbations (position and orientation jointly) on map quality in Forest II and III (Fig. 5 (c) and (d), respectively). As observed, the average KLD (all voxels) on the expected occupancy also outperforms the standard occupancy with combined perturbation values.
Finally, we computed the Area Under Curve (AUC) of the Receiver Operating Characteristic (ROC) curves for the standard occupancy and expected occupancy approaches, providing an overview of each method’s performance as predictors of free or occupied states in datasets Forest II and III. Results are summarised in Table I, where AUC values closer to one indicate better prediction rates. For increasing levels of perturbation, both approaches experience reduction of AUC, which is anticipated since higher errors on sensor poses negatively affect the quality of occupancy values in both methods. However, in both datasets, the expected occupancy method displayed higher AUC than the standard occupancy, showing the benefits of incorporating localisation uncertainty in the occupancy calculation.
TABLE I: Comparison of the AUC values for the standard occupancy (occ) and expected occupancy (exp) approaches for increasing perturbation levels in sensor poses, where $\sigma=(0.015\text{m},0.015\text{m},0.015\text{m},0.11^{\circ},0.11^{\circ},0.11^{\circ})$. Higher AUC values indicate better map predictions.
Forest II Forest III
Perturbation level occ exp occ exp
$\sigma$ 0.982 0.988 0.984 0.990
[…]
VI-B Full system evaluation
To assess whether our navigation approach (mapping and planning), produces more efficient paths in comparison with a naive baseline planner without prior maps, we also created a simulated forest with ground obstacles as illustrated in Fig. 6. Similarly to Sec. VI-A, we generated a lidar trajectory above the canopy with poses perturbed using $\sigma=(0.05\text{m},0.05\text{m},0.05\text{m},0.11^{\circ},0.11^{\circ},0.11^{\circ})$. We then processed the lidar scans using our mapping method (Sec. IV) to create the obstruction map used by the D* Lite planner.
Figure 6: (a) Gazebo forest model used in simulations of the planning methods. (b) Shows the density of ground obstacles (red) composed by high vegetation and tree trunks in the synthetic forest.
Figure 7: Executed paths comparison between our planner using three cost functions: log-reachability cost, expectation cost with ($C_{\text{obst}}:5$) and ($C_{\text{obst}}:20$), against the naive planner using only local information. (a) shows the histogram of path length ratio between the executed path of our method and the naive approach. (b) depicts the cumulative count of paths for the same ratios. Curves closer to the top left show better performance.
Figure 8: The satellite image (a) and the aggregated lidar scans (b) of the forest area used in our experiments. The resulting obstruction map generated using our proposed approach is shown in (c) along with a magnified section for better visualisation in (d).
We defined 40 random start and end goal pairs in the synthetic forest to be executed by our navigation method and the naive planner. The naive planner does not use the prior map, and is equivalent to assuming a direct path from the start to the goal as the global prior path. All 40 runs were repeated with three configurations of cost function: log-reachability cost, and two versions of expected cost (5) with values $C_{\text{obst}}=5$ and $C_{\text{obst}}=20$, respectively. During online execution of the planners, the ground-truth is utilised to update the obstruction map in our navigation approach and to construct a map on the fly in the naive baseline. We defined a 5 meter radius around the robot pose to update the maps with local information from the ground-truth environment, hence map cells in this region are changed to near 1 or 0, in case obstacles or free areas are detected, respectively, in the same cell coordinates of the ground-truth map. In both methods we utilise the D* Lite planner to replan in case obstacles block the robot’s planned path. While our method replans on a combined map (prior map with local updates), the naive approach replans only on the map built online with local information.
The metric used for evaluation is the ratio of the executed path length between our navigation method and the naive baseline. Fig. 7 (a) and (b) show the results of our complete navigation framework using the chosen configurations of cost function. As observed in the histogram, using the log-reachability cost resulted in $87.5\%$ of executed paths having lengths shorter than the naive planner, confirming the clear advantage of using our mapping method as prior information for global planning and also online execution and replanning. This result also confirms that even in the presence of perturbations in the sensor trajectory and partial canopy cover, the proposed expected occupancy and obstruction mapping approaches provide enough information for generating efficient paths.
The expectation cost, however, resulted in $47.5\%$ of paths shorter than the baseline. This performance reduction can be attributed to the fact that when using the expected cost some planned paths crossed a small number of cells with high obstruction scores instead of planning longer detours. During execution with the ground robot, some of these paths tended to guide the robot to blocked areas in the ground-truth map (in some instances multiple times), requiring expensive local replanning steps. Further, although the naive approach has no prior, in cases with few obstructions the naive assumption (straight-line) may actually be globally optimal or near-optimal, requiring little or no replanning. Because the log-reachability cost is more risk averse by design, it may prefer to detour even small potentially obstructed areas, yielding better results (see Fig. 7 (b)).
VII Real-World Experiments
We also evaluated our offline mapping and planning framework using real aerial view lidar datasets and a ground mobile robot. The airborne lidar scans were collected using a DJI M-300 quadcopter equipped with a velodyne VLP-16 lidar sensor. For ground deployment, we used a Dynamic Tracked Robot (DTR) (shown in Fig.1) equipped with our in-house lidar perception pack[24] for local sensing and obstacle avoidance. The tests were performed in the western forest area at the Queensland Centre for Advanced Technologies (QCAT) in Brisbane, Australia. The testing area has dense canopy cover and varied obstacles over the terrain, e.g., high vegetation, logs, trees and narrow gaps, which makes the global navigation and local traversability highly complex.
Figure 9: Results of real world experiments Test 1 (a) and Test 2 (b) showing the executed paths of our approach and the naive method. Black circle and star represent the start and goal locations, respectively. In both experiments our approach guided the robot to shorter and less cluttered paths, whereas the limited look-ahead of the naive method resulted in a longer trajectory in (a) due to crossing a more cluttered area, and ultimate failure of the planner in (b) by facing a dead-end of high undergrowth.
Since it is not trivial to define ground-truth for obstructed areas in forest environments given their complexity, e.g., cells that are visually “obstructed” by vegetation are often traversable, we analysed qualitatively our mapping output from real data. As observed in Fig. 8 (a)-(d), even with partial canopy occlusion, our mapping using above-canopy lidar scans is capable of detecting regions at ground level with high obstruction scores (yellow to red) indicating the presence of obstacles, and free regions (green) with higher chances of traversing. Uncertain areas (blue) are also captured, providing meaningful information for planning and decision-making.
To test our full system, we defined two start and end goal pairs in the forest named Test 1 and 2. For each test location we performed two runs, one with the naive approach where the robot used only the local perception and its embedded navigation system to reach the target without any prior information, and another run using our proposed navigation method (prior aerial mapping and global planning) with the log-reachability cost as it demonstrated superior performance in simulation. Unlike the navigation pipeline used in simulations, here our method was used only offline to create the obstruction map and plan a global trajectory from start to goal using D* Lite. Then, for online execution, the only prior information provided to the robot was the planned trajectory as a set of waypoints followed by the local navigation system [23] of the DTR. The local planner (Hybrid A* [30]) and perception allow the robot to deviate around locally-detected obstacles while tracking the waypoints. For fair comparison, the same local navigation system was used for both runs, naive and our method; in this manner, we can assess the benefits of providing information prior to operation through a global trajectory.
During Test 1, depicted in Fig. 9 (a), while our approach preferred a less cluttered path on the left side of the large obstacle, the naive approach took the right side, following a much denser path that also crossed a narrow passage among high bushes. Our method travelled a path length of $69.6$ m in comparison to $79.9$ m of the naive trajectory, which demonstrates that our proposed navigation not only increases path efficiency, but it is also more risk averse preferring safer areas with less likelihood of having obstacles.
In Test 2, Fig. 9 (b), the limited look-ahead of the naive approach directed the robot to a dead-end composed of cluttered high vegetation. The local navigation was unable to detour the blocked area, and failed to reach the goal. Unlike the naive counterpart, our method leveraged the prior information, avoiding the dead-end and taking a clear path on the right of the vegetated area to reach the goal.
Although we acknowledge that the practical experiments do not represent statistically representative results, the outcomes validate the beneficial behaviours of our method observed in simulation. The field trials showed that planning using our obstruction maps in tandem with the proposed log-reachability cost leads to shorter and safer paths by avoiding cluttered areas, which ultimately increases the mission success rate.
VIII Conclusion & Future Work
We presented a navigation system for mapping using prior overhead lidar data and planning in complex forest environments. Our proposed mapping estimates the obstructed areas on ground level using a scoring function of occupancy values obtained from airborne lidar scans and the sensor trajectory. A Monte Carlo sampling approach to incorporate sensor pose uncertainty in occupancy maps is also presented and extensively tested in simulation, showing more accurate maps than the standard occupancy mapping. We proposed two path cost functions that take into account both obstruction uncertainty and path length. Lastly, D* Lite leverages the cost functions to plan optimal paths for ground-navigation, also allowing efficient replanning. Simulations and real experiments demonstrated that our navigation approach using prior information leads to shorter paths than the naive baseline (no prior) and increases the chances of reaching the goal even in challenging forests scenarios. In future work we intend to learn the obstruction score map in a self-supervised fashion online, improving the generalisation and autonomy of our approach.